HETAV PATEL

Grade 11,

Sir, Please give a note on Arrangements of BCC & FCC Thanks

Sir, Please give a note on Arrangements of BCC & FCC
 
Thanks

Grade:12th pass

1 Answers

Mahima Kanawat
5 years ago
Dear student 
BCC is body centred cubic. The atoms are present on all the corners and body centre.
FCC is face centred cubic. The atoms are present on all the corners and face centre. 
Number of atoms in Lattice can be calculated by their contributions. 
For corner atoms contribution is 1/8
For body centre atom 1
For face centre atom 1/2
So, in bcc, total atoms = 8*1/8+1*1=2
For fcc, total atoms = 8*1/8+6*1/2 = 4
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT